The formula h=50t-5t(squared)can be used to find the height in meters of an object shot upward at an initial velocity of 50 meters per second after t seconds. Find the time that the object will be 120 meters high on its way up
Answer by Earlsdon(6294)

Here, you are given the height (h) and are asked to find the time (t) at which this height will occur. The problem is solved by substituting the given value of h (120 meters) into the formula, and solving for t.
120 = 50t - 5t^2 Rewrite this into standard quadratic equation form:
-5t^2 + 50t - 120 = 0 Divide through by -5 to facilitate the calculation.
t^2 - 10t + 24 = 0 Solve this quadratic equation by factoring.
(t - 12)(t + 2) = 0 Apply the zero products principle.
t - 12 = 0, t = 12
t + 2 = 0, t = -2 Discard this solution as a negative value for time is not meaningful.
Answer: 12 seconds.